OpenStudy (anonymous):

Jose is rappelling down a 526-foot cliff. He moves down the cliff at a rate of 8 feet per second. Which equation would represent the height Jose is above the ground, H, after t seconds?

OpenStudy (anonymous):

is there options

OpenStudy (anonymous):

A. H = 526 - 8t B. H = 526 + 8t C. H = 8 - 526t D. H = 8 + 526t

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

lol are you guessing?

OpenStudy (anonymous):

I would have to also go with A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks